Question 15 Waterway Industries purchased equipment in January of 2008 for $402000. The equipment was being depreciated on the straight-line method over an estimated useful life of 20 years, with no salvage value. At the beginning of 2018, when the equipment had been in use for 10 years, the company paid $51000 to overhaul the equipment. As a result of this improvement, the company estimated that the useful life of the equipment would be extended an additional 5 years. What should be the depreciation expense recorded for this equipment in 2018? $16800o O $20100 O $13350 O $10050
Answer #1

Yearly depreciation charged from January 2008 to December 2017(10 years)= (402000/20)*10= $201000

Formula used to compute depreciation under straight line method= (purchase cost of asset-salvage value)/life of the asset

Value of the asset as on 1 January 2018=Purchase cot- depreciation charged till dec 2017

=402000-201000=$201000

Cost of improvement= $51000 (to be included in the value of the asset as it increases the life of the asset ,ie, the production capacity of the asset)

Value of asset after improvement= 201000+51000= $252000

Depreciation to be charged on 2018 and thereon every year= 252000/(10+5)= $16800

Option a) is the correct answer.